3. When iron reacts with thiocyanate, a deep red color complex is formed. Which wavelength do you think it will show maximum absorbance (λmax).

4. It has been experimentally shown that the absorbance of a given Fe(III)/thiocyanate solution decreases as the temperature increases. Can you think of some possible explanations for this behavior?
Answer #1

#3. Colors of complexes absorbs light as per VIBGYOR from ;

Color = wavelength of light at which maximum absorbance.

Violet = 400 - 440 nm

Indigo = 440 - 460 nm

Blue = 460 - 500 nm

Green = 500 - 570 nm

Orange = 570 - 620

Red = 620 - 740 nm

Now as the complex formed is of deep red color , then it will show maximum absorbance at ~ 680 nm wavelength.

#4. As the temperature is raised , the complex dissociates into the consituting ions ( Fe^+3 and SCN^- ions ) and thus deep red color's intensity is decreased with increasing temperature , that is why wavelength of maximum absorbance decreased .